An observer sees that the angle of elevation to the top of a tower, 80m high and due south of her position is 37 degrees. The top of another tower 60 m high and due west of her position is at an angle of elevation of 28 degrees. Calculate the distance between the two towers.
Answer by mananth(16949)

ground distance of the south tower = 80/Tan 37
ground distance of west tower = 60/Tan 27

The plot of towers and the point of observation form a right triangle on ground
By Pythagoras theorem

distance between towers = sqrt ((60/tan 27 degrees)^2 +80/tan 37 degrees)^2)
= 158.5 meters